# A contextual advantage for conclusive exclusion: repurposing the Pusey-Barrett-Rudolph construction

### v1

@@ Appendix B Proof of the noncontextuality inequality
$\beta_{0-}=\int_{\Lambda_{0-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{-}\}d\lambda$
$\beta_{1+}=\int_{\Lambda_{1+}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{+}\}d\lambda$
$\beta_{1-}=\int_{\Lambda_{1-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{-}\}d\lambda,$
Let us begin by focusing on the first term in each $\beta_{T}$. Taking the term $\int_{\Lambda_{0+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda$ in $\beta_{0+}$ as an example. For any $\lambda\in\Lambda_{0+}$, we have $\mu_{1}(\lambda)=\mu_{-}(\lambda)=0$. Together with Eq. 16, this implies that $\mu_{0}(\lambda)=\mu_{+}(\lambda)$. Thus $\forall\lambda\in\Lambda_{0+}$, we have $\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}=\mu_{0}(\lambda)$ or $\mu_{+}(\lambda)$. Hence,
@@
$\forall\lambda\in\Lambda_{0-},\,\mu_{0}(\lambda)=\mu_{-}(\lambda);\quad\forall\lambda\in\Lambda_{1+},\,\mu_{1}(\lambda)=\mu_{+}(\lambda);\quad\forall\lambda\in\Lambda_{1-},\,\mu_{1}(\lambda)=\mu_{-}(\lambda).$
@@
$\beta_{0+}=\int_{\Lambda_{0+}}\mu_{0}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{+}\}d\lambda$
$\beta_{0-}=\int_{\Lambda_{0-}}\mu_{0}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{-}\}d\lambda$
$\beta_{1+}=\int_{\Lambda_{1+}}\mu_{1}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{+}\}d\lambda$
$\beta_{1-}=\int_{\Lambda_{1-}}\mu_{1}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{-}\}d\lambda,$
@@
$\int_{\Lambda_{0+-}}[\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}+\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}]d\lambda=$ $\int_{\Lambda_{0+-}}[\mu_{+}(\lambda)+\mu_{-}(\lambda)]d\lambda$
$=$ $\int_{\Lambda_{0+-}}[\mu_{0}(\lambda)+\mu_{1}(\lambda)]d\lambda,$
@@
${\rm CE}=3.$ (41)
This experiment admits of a noncontextual ontological model since all states and measurements are in the stabilizer subtheory of the qubit theory, and it is well known than any prepare-measure experiment of the stabilizer qubit subtheory admits of a noncontextual ontological model [Lillystone2019].

### v2

@@ Appendix B Proof of the noncontextuality inequality
$\beta_{0-}=\int_{\Lambda_{0-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{-}\}d\lambda,$ (37)
$\beta_{1+}=\int_{\Lambda_{1+}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{+}\}d\lambda,$ (38)
$\beta_{1-}=\int_{\Lambda_{1-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{-}\}d\lambda.$ (39)
Let us begin by focusing on the first term in each $\beta_{T}$. Taking the term $\int_{\Lambda_{0+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda$ in $\beta_{0+}$ as an example. For any $\lambda\in\Lambda_{0+}$, we have $\mu_{1}(\lambda)=\mu_{-}(\lambda)=0$. Together with Eq. 16, this implies that $\mu_{0}(\lambda)=\mu_{+}(\lambda)$. Thus $\forall\lambda\in\Lambda_{0+}$, we have $\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}=\mu_{0}(\lambda)=\mu_{+}(\lambda)$. Hence,
@@
$\forall\lambda\in\Lambda_{0-},\,\mu_{0}(\lambda)=\mu_{-}(\lambda);\quad\forall\lambda\in\Lambda_{1+},\,\mu_{1}(\lambda)=\mu_{+}(\lambda);\quad\forall\lambda\in\Lambda_{1-},\,\mu_{1}(\lambda)=\mu_{-}(\lambda).$ (41)
@@
$\beta_{0+}=\int_{\Lambda_{0+}}\mu_{0}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{+}\}d\lambda,$ (42)
$\beta_{0-}=\int_{\Lambda_{0-}}\mu_{0}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{0},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{-}\}d\lambda,$ (43)
$\beta_{1+}=\int_{\Lambda_{1+}}\mu_{1}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{1},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{+}\}d\lambda,$ (44)
$\beta_{1-}=\int_{\Lambda_{1-}}\mu_{1}d\lambda+\int_{\Lambda_{1+-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01-}}\min\{\mu_{1},\mu_{-}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{1},\mu_{-}\}d\lambda.$ (45)
@@
$\int_{\Lambda_{0+-}}[\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}+\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}]d\lambda=\int_{\Lambda_{0+-}}[\mu_{+}(\lambda)+\mu_{-}(\lambda)]d\lambda=\int_{\Lambda_{0+-}}[\mu_{0}(\lambda)+\mu_{1}(\lambda)]d\lambda,$ (46)
@@
${\rm CE}=3.75.$ (55)
This experiment admits of a noncontextual ontological model since all states and measurements are in the stabilizer subtheory of the qubit theory, and it is well known that any prepare-measure experiment of the stabilizer qubit subtheory admits of a noncontextual ontological model [Lillystone2019].